# Detecting Toxic Flow

### v1

@@ Abstract
This paper develops a framework to predict toxic trades that a broker receives from her clients. Toxic trades are predicted with a novel online Bayesian method which we call the projection-based unification of last-layer and subspace estimation (PULSE). PULSE is a fast and statistically-efficient online procedure to train a Bayesian neural network sequentially. We employ a proprietary dataset of foreign exchange transactions to test our methodology. PULSE outperforms standard machine learning and statistical methods when predicting if a trade will be toxic; the benchmark methods are logistic regression, random forests, and a recursively-updated maximum-likelihood estimator. We devise a strategy for the broker who uses toxicity predictions to internalise or to externalise each trade received from her clients. Our methodology can be implemented in real-time because it takes less than one millisecond to update parameters and make a prediction. Compared with the benchmarks, PULSE attains the highest PnL and the largest avoided loss for the horizons we consider.
@@ 1 Introduction
Liquidity providers are key to well-functioning financial markets. In foreign exchange (FX), as in other asset classes, broker-client relationships are ubiquitous. The broker streams bid and ask quotes to her clients and the clients decide when to trade on these quotes, so the broker bears the risk of adverse selection when trading with better informed clients. These risks are borne by both liquidity providers who stream quotes to individual parties and by market participants who provide liquidity in the books of electronic exchanges. However, in contrast to electronic order books in which trading is anonymous for all participants (e.g., in Nasdaq, LSE, Euronext), in broker-client relationships the broker knows which client executed the order. This privileged information can be used by the broker to classify flow, i.e., toxic or benign, and to devise strategies that mitigate adverse selection costs.
@@
In the literature, models generally classify traders as informed or uninformed; see e.g., Bagehot, (1971), Copeland and Galai, (1983), Grossman and Stiglitz, (1980), Amihud and Mendelson, (1980), Kyle, (1989), Kyle, (1985), and Glosten and Milgrom, (1985). In equity markets, many studies focus on informed flow (i.e., asymmetry of information) across various traded stocks, see e.g., Easley et al., (1996) who study the probability of informed trading at the stock level, while our study focuses on each trade because we have trader identification. In FX markets, Butz and Oomen, (2019) develop a model for internalisation in FX markets, and Oomen, (2017) studies execution in an FX aggregator and the market impact of internalisation-externalisation strategies. Overall, studies of toxic flow and information asymmetry do not make predictions of toxicity at the trade level. To the best of our knowledge, ours is the first paper in the literature to use FX data with trader identification to predict the toxicity of each trade.
@@
In our work, a trade is toxic if a client can unwind the trade within a given time window and make a profit (i.e., a loss for the broker). Toxic trades are not necessarily informed, nor informed trades are necessarily toxic. An uninformed client can execute a trade that becomes toxic for the broker because of the random fluctuations of exchange rates. Ultimately, the broker’s objective is to avoid holding loss-leading trades in her books, so it is more effective to focus on market features and each trade the broker fills, rather than on whether a particular client is classified as informed or uninformed. For simplicity, theoretical models in the literature assume traders are informed or uninformed, while in practice not all trades sent by one particular client are motivated by superior information.111 Cartea and Sánchez-Betancourt, (2022) show how the broker optimally provides liquidity to clients who have been previously classified as informed or uninformed.
@@
The main contributions of our paper are as follows. We predict the toxicity of each incoming trade with machine learning and statistical methods, such as logistic regression, random forests, a recursively updated maximum-likelihood estimator, and develop a novel algorithm that uses a neural network (NNet) and that is able to update the model parameters sequentially and efficiently. We call this new method PULSE, which stands for projection-based unification of last-layer and subspace estimation.
@@
Our new method employs a neural network to compute the probability that a trade will be toxic. After revealing if a trade was toxic, PULSE uses the last observed features to update the parameters of the NNet. To estimate parameters efficiently at each timestep, we follow three steps. One, we split the last layer from the feature-transformation layers of a NNet. Two, we project the parameters of the feature-transformation layers onto an affine subspace. Three, we devise a recursive formula to estimate a posterior distribution over the projected feature-transformation parameters and last-layer parameters. Specifically, we extend the subspace neural network model (subspace NNets) of Duran-Martin et al., (2022) to classification tasks. We also use the exponential-family extended Kalman filter (expfam EKF) method of Ollivier, (2017) and we follow the ideas of the recursive variational Gaussian approximation (R-VGA) results of Lambert et al., (2021) to obtain the update equations in PULSE. Finally, we impose a prior independence between the hidden layers of the neural network and the output layer, extending the work in the last-layer Bayesian neural networks literature (last-layer BNNs). Figure 1 shows the relationship of PULSE to previous methods.
@@
To evaluate the predictive performance of our model and the efficacy of the broker’s strategy, we use a proprietary dataset of FX transactions from 28 June 2022 to 21 October 2022. Initially, the models are trained with data between 28 June and 31 July, and the remainder of the data (1 August to 21 October) is used to deploy the strategy, i.e., use predictions of toxicity for each trade to inform the internalisation-externalisation strategy we develop. During the deploy phase, PULSE and the maximum-likelihood estimator models continue to learn, while the models based on logistic regression and random forests are not updated. For a given toxicity horizon and a cutoff probability the (internalisation-externalisation) strategy internalises the trade if the probability that the trade is toxic is less than or equal to the cutoff probability, otherwise it externalises the trade. We compute the PnL of all trades that the broker internalised and the avoided loss due to externalised trades. We find that PULSE delivers the best combination of PnL and avoided loss across all toxicity horizons we consider in this paper.
@@
Finally, it is more advantageous to have a universal model than to have one model per trader. That is, we obtain higher accuracies when we train one model for all traders than when we train one model per trader; higher accuracies result from having more data. When one restricts to one model per trader, the model for traders with fewer transactions underperforms compared with a universal model that is trained on more datapoints. We also find that if we build a model that does not consider the inventory, cash, and recent activity of clients, that is, if the broker does not use identification of the trader, the performance of PULSE is the same. Thus, in our dataset, client-specific variables do not add value to predict the toxicity of trades.
2 Data analysis
2.1 The data
@@
We employ data for the currency pair EUR/USD from LMAX Broker and from LMAX Exchange for the period 28 June 2022 to 21 October 2022.222www.lmax.com. For each liquidity taking trade filled by the broker, we use the direction of trade (buy or sell), the timestamp when LMAX Broker processed the trade, and the volume of the trade. Also, we use the best quotes and volumes available in LMAX Exchange at a microsecond frequency. In contrast to LMAX Broker, traders who interact in the limit order book (LOB) of LMAX Exchange do not know the identity of their counterparties. The LOB uses a price-time priority to clear supply and demand of liquidity — as in traditional electronic order books in equity markets, such as those of Nasdaq, Euronext, and the London Stock Exchange.
@@
Below, we work with the data of Clients 1 to 6 in Table 1 and we assume that the broker quotes her clients the best available rates in LMAX Exchange net of fees. Transaction costs in FX are around \$3 per million euros traded (see e.g., Cartea and Sánchez-Betancourt, (2023)), so this assumption provides clients with a discount of \$3 per million euros traded when trading with the broker.
2.2 Toxicity
@@
Denote time by $t\in{\mathfrak{T}}=[0,T]$, where 0 is the start of the trading day and $T$ is the end of the trading day. From this point forward, we use ‘exchange rate’ and ‘prices’ interchangeably. The best ask price and best bid price in the LOB of LMAX Exchange are denoted by $(S^{a}_{t})_{t\in{\mathfrak{T}}}$ and $(S^{b}_{t})_{t\in{\mathfrak{T}}}$, respectively. Let $\mathfrak{G}$ be a toxicity horizon such that $0<\mathfrak{G}\ll T$, and let $t\in[0,T-\mathfrak{G}]$. We define the two stopping times
@@
with the convention that $\inf\,\emptyset=\infty$. The stopping time $\tau^{+}_{t}$ is the first time after $t$ that the best bid price is above the best ask price at time $t$. If $\tau^{+}_{t}<\infty$, a buy trade executed at $S_{t}^{a}$ becomes profitable for the client at time $\tau^{+}_{t}$ before the end of the trading day because the client can unwind her position and collect the profit
@@
Let $\mathfrak{G}>0$ be a toxicity horizon. A client’s buy (resp. sell) filled by the broker at time $t$ is toxic for the broker if $\tau^{+}_{t}\leq t+\mathfrak{G}$ (resp. if $\tau^{-}_{t}\leq t+\mathfrak{G}$).
@@
Figure 2: A sell trade from a client that becomes toxic for the broker after a few seconds of filling the trade.
In the figure, if a client buys from the broker at the best ask price at time $t=$10:00:00 am, then there is no opportunity for the trader to unwind the trade at a profit in the first ten seconds after the trade. However, had the trader sold to the broker at the best bid price at time $t=$10:00:00 am, then shortly after 10:00:01 the trade would be in-the-money for the client (i.e., toxic for the broker).
@@
toxicity horizon $\mathfrak{G}$ in seconds
@@ 2.3 Features to predict toxic trades
[…]
$\text{sign}\left(\mathfrak{Q}^{c}_{t^{-}}\right)\times\log\left(1+|\mathfrak{Q}^{c}_{t^{-}}|\right)\,,$
[…]
$\mathfrak{Q}^{c}_{t}=\int_{0}^{t}q^{c}_{u}\,\mathrm{d}N^{c,a}_{u}-\int_{0}^{t}q^{c}_{u}\,\mathrm{d}N^{c,b}_{u}\,.$
Here, $q^{c}_{t}$ is the size of the order sent at time $t$ by client $c$ and $N^{c,a}_{t}$, $N^{c,b}_{t}$ are the counting processes of buy and sell orders, respectively, sent by client $c$ and filled by the broker. The cash of client $c$, denoted by $\mathfrak{C}^{c}_{t}$, is given by
$\mathfrak{C}^{c}_{t}=-\int_{0}^{t}S^{a}_{u^{-}}\,q^{c}_{u}\,\mathrm{d}N^{a,c}_{u}+\int_{0}^{t}S^{b}_{u^{-}}\,q^{c}_{u}\,\mathrm{d}N^{b,c}_{u}\,,$
and the feature associated with the cash process is
$\text{sign}\left(\mathfrak{C}^{c}_{t^{-}}\right)\times\log\left(1+|\mathfrak{C}^{c}_{t^{-}}|\right)\,.$
In LMAX Exchange, the bid-ask spread is
$\mathfrak{S}_{t}=S^{a}_{t}-S^{b}_{t}\,,$
the midprice is
$S_{t}=\frac{S^{a}_{t}+S^{b}_{t}}{2}\,,$
the volume imbalance of the best available volumes is defined by
$\mathfrak{I}_{t}=\frac{V_{t}^{\text{b}}-V_{t}^{\text{a}}}{V_{t}^{\text{b}}+V_{t}^{\text{a}}}\,,$
and the associated feature for the volume $V$ is the transformed volume
$\mathfrak{V}=\log(1+|V|)\,.$
The number of trades received by the broker from her clients is
$N_{t}=\sum_{c\in{\mathcal{A}}}\left(N^{c,a}_{t}+N^{c,b}_{t}\right)\,,$
and the volatility of the midprice in the LOB of LMAX Exchange over the interval $[t-\delta,t)$ is the square root of the quadratic variation of the logarithm of the midprice over the interval. More precisely,
$\mathfrak{V}^{\delta}_{t}=\sqrt{\sum_{\Delta\log{S_{u}}\neq 0\,;\,u\in[t-\delta,t)}|{\Delta\log{S_{u}}}|^{2}}\,,$
where
$\log\left({S_{t^{-}}}/{S_{t-\delta}}\right).$
The timing of the events in the LOB is measured with three clocks: time-clock, transaction-clock, and volume-clock. The time-clock runs as $t\in[0,T]$ with microsecond precision, i.e., a millionth of a second. For a given day $d$ with $N^{d}$ transactions and $V^{d}$ volume traded, the transaction clock runs as $\mathfrak{t}\in[0,N^{d}]$, and the volume-clock runs as $\mathfrak{v}\in[0,V^{d}]$. The number of transactions $\mathfrak{t}$ is a function of $t$, that is, $\mathfrak{t}(t)$ is the number of transactions observed up until time $t$, similar for the volume-clock $\mathfrak{v}(t)$. Thus, for any order sent at time $t$, the time associated with the order in the transaction-clock is $\mathfrak{t}(t)$ and the time in the volume-clock is $\mathfrak{v}(t)$.
@@
For each clock and a given interval in the past (as measured by the clock), we employ the following eight features: (a) volatility of midprice, (b) number of trades executed by the client in the interval, (c) number of updates in the best quotes of LMAX Exchange in the interval, (d) return of midprice over the interval, (e) average transformed volume in the best bid price, (f) average transformed volume in the best ask price, (g) average spread, and (h) average imbalance of the best available volumes. Here, average refers to sample average and not clock-weighted average; our results are similar with either average. We use sample averages because it speeds up the deployment of the strategy.
[…]

### v2

@@ Abstract
This paper develops a framework to predict toxic trades that a broker receives from her clients. Toxic trades are predicted with a novel online learning Bayesian method which we call the projection-based unification of last-layer and subspace estimation (PULSE). PULSE is a fast and statistically-efficient Bayesian procedure for online training of neural networks. We employ a proprietary dataset of foreign exchange transactions to test our methodology. Neural networks trained with PULSE outperform standard machine learning and statistical methods when predicting if a trade will be toxic; the benchmark methods are logistic regression, random forests, and a recursively-updated maximum-likelihood estimator. We devise a strategy for the broker who uses toxicity predictions to internalise or to externalise each trade received from her clients. Our methodology can be implemented in real-time because it takes less than one millisecond to update parameters and make a prediction. Compared with the benchmarks, online learning of a neural network with PULSE attains the highest PnL and avoids the most losses by externalising toxic trades.
@@ 1 Introduction
Liquidity providers are key to well-functioning financial markets. In foreign exchange (FX), as in other asset classes, broker-client relationships are ubiquitous. The broker streams bid and ask quotes to her clients and the clients decide when to trade on these quotes, so the broker bears the risk of adverse selection when trading with better informed clients. These risks are borne by both liquidity providers who stream quotes to individual parties and by market participants who provide liquidity in the books of electronic exchanges. However, in contrast to electronic order books in which trading is anonymous for all participants (e.g., Nasdaq, LSE, Euronext), in broker-client relationships the broker knows which client executed the order. This privileged information can be used by the broker to classify flow, i.e., toxic or benign, and to devise strategies that mitigate adverse selection costs.
@@
In the literature, models generally classify traders as informed or uninformed; see e.g., Bagehot (1971), Copeland and Galai (1983), Grossman and Stiglitz (1980), Amihud and Mendelson (1980), Kyle (1989), Kyle (1985), and Glosten and Milgrom (1985). In equity markets, many studies focus on informed flow (i.e., asymmetry of information) across various traded stocks, see e.g., Easley et al. (1996) who study the probability of informed trading at the stock level, while our study focuses on each trade because we have trader identification. In FX markets, Butz and Oomen (2019) develop a model for internalisation, and Oomen (2017) studies execution in an FX aggregator and the market impact of internalisation-externalisation strategies. Overall, studies of toxic flow and information asymmetry do not make predictions of toxicity at the trade level. To the best of our knowledge, ours is the first paper in the literature to use FX data with trader identification to predict the toxicity of each trade.
@@
In our work, a trade is toxic if a client can unwind the trade within a given time window and make a profit (i.e., a loss for the broker). Toxic trades are not necessarily informed, nor informed trades are necessarily toxic. An uninformed client can execute a trade that becomes toxic for the broker because of the random fluctuations of exchange rates. Ultimately, the broker’s objective is to avoid holding loss-leading trades in her books, so it is more effective to focus on market features and on each trade the broker fills, rather than on whether a particular client is classified as informed or uninformed. For simplicity, theoretical models in the literature assume traders are informed or uninformed, while in practice not all trades sent by one particular client are motivated by superior information.
@@
The main contributions of our paper are as follows. We predict the toxicity of each incoming trade with machine learning and statistical methods, such as logistic regression, random forests, a recursively updated maximum-likelihood estimator, and a neural network (NNet). We devise a novel algorithm to update the parameters of the NNet sequentially; we call this rule of learning PULSE, which stands for projection-based unification of last-layer and subspace estimation. We deploy our toxicity prediction models in a proprietary dataset, and we find that using a single model for all clients (employing client-specific features) outperforms the use of one model per client. We also find that, compared with the benchmarks, the methodology we put forward attains the highest PnL and avoids the most losses by externalising toxic trades.
@@
Our new method employs a NNet to compute the probability that a trade will be toxic. After the outcome of each trade, toxic or benign, PULSE updates the parameters of the NNet. To update the parameters efficiently at each timestep, PULSE follows three steps: one, split the last layer from the feature-transformation layers of a NNet; two, project the parameters of the feature-transformation layers onto an affine subspace; and three, devise a recursive formula to estimate a posterior distribution over the projected feature-transformation parameters and last-layer parameters. Specifically, we extend the subspace NNet model (subspace NNets) of Duran-Martin et al. (2022) to classification tasks. We also use the exponential-family extended Kalman filter (expfam EKF) method of Ollivier (2017) and we follow the ideas of the recursive variational Gaussian approximation (R-VGA) results of Lambert et al. (2021) to obtain the update equations in PULSE. Finally, we impose a prior independence between the hidden layers of the NNet and the output layer, extending the work in the last-layer Bayesian NNet (last-layer BNNs). Figure 1 shows the relationship of PULSE to previous methods. In short, PULSE is a statistically-efficient update rule to learn the parameters of a NNet sequentially.
@@
To evaluate the predictive performance of our model and the efficacy of the broker’s strategy, we use a proprietary dataset of FX transactions from 28 June 2022 to 21 October 2022. Initially, the models are trained with data between 28 June and 31 July, and the remainder of the data (1 August to 21 October) is used to deploy the strategy, i.e., use predictions of toxicity for each trade to inform the internalisation-externalisation strategy we develop. During the deploy phase, the maximum-likelihood estimator is continually updated with the running average of the toxic trades and the parameters of the NNet are updated with PULSE, while the models based on logistic regression and random forests are not updated. [^1] Robustness checks in the appendix study the performance of the logistic regression and random forests when the models are re-trained weekly. For a given toxicity horizon and a cutoff probability, the strategy internalises the trade if the probability that the trade is toxic is less than or equal to the cutoff probability, otherwise it externalises the trade. We compute the PnL of all trades that the broker internalised and the losses she avoids by externalising trades. We find that a NNet trained with PULSE delivers the best combination of PnL and avoided loss across all toxicity horizons we consider in this paper.
@@
Finally, we find that a universal model is more advantageous than one model per trader. That is, we obtain higher accuracies (when predicting toxic trades) when we train one model for all traders than when we train one model per trader; higher accuracies result from having more data. When one restricts to one model per trader, the model for traders with fewer transactions is outperformed by a universal model that is trained on more datapoints. We also find that if we build a universal model that does not consider the inventory, cash, and recent activity of clients (i.e., the broker does not use the identification of the trader), the performance of PULSE deteriorates substantially when compared to the performance of a model that includes the identity and unique features of the trader. Thus, in our dataset, client-specific variables does add value to predict the toxicity of trades.
The remainder of paper is organised as follows. Section 2 describes the data. Section 3 defines toxicity, provides statistics about the clients in the dataset, and illustrates the toxicity profiles of clients. Section 4 introduces PULSE, which is a fast and statistically-efficient Bayesian procedure for online training of neural networks. Section 5 shows implementation details. Section 6 uses proprietary datasets to evaluate PULSE against alternative methods. Section 7 presents conclusions. We collect proofs, together with additional robustness checks, in the appendix.
2 Data and preliminary analysis
@@
We employ data for the currency pair EUR/USD from LMAX Broker and from LMAX Exchange for the period 28 June 2022 to 21 October 2022.222www.lmax.com. For each liquidity taking trade filled by the broker, we use the direction of trade (buy or sell), the timestamp when LMAX Broker processed the trade, and the volume of the trade. Also, we use the best quotes and volumes available in LMAX Exchange at a microsecond frequency. In contrast to LMAX Broker, traders who interact in the limit order book (LOB) of LMAX Exchange do not know the identity of their counterparties. The LOB uses price-time priority to clear supply and demand of liquidity — as in traditional electronic order books in equity markets, such as those of Nasdaq, Euronext, and the London Stock Exchange.
@@
Below, we work with the data of Clients 1 to 6 and we assume that the broker quotes her clients the best available rates in LMAX Exchange net of fees. Transaction costs in FX are around \$3 per million euros traded (see e.g., Cartea and Sánchez-Betancourt (2023)), so this assumption provides clients with a discount of \$3 per million euros traded when trading with the broker. [^3] In practice, not all clients receive the same bid and ask quotes.
3 Toxicity
@@
Denote time by $t\in{\mathfrak{T}}=[0,T]$, where 0 is the start of the trading day and $T$ is the end of the trading day. From this point forward, we use ‘exchange rate’ and ‘prices’ interchangeably. The best ask price and best bid price in the LOB of LMAX Exchange are denoted by $(S^{a}_{t})_{t\in{\mathfrak{T}}}$ and $(S^{b}_{t})_{t\in{\mathfrak{T}}}$, respectively. Let $\mathcal{G}$ be a toxicity horizon such that $0<\mathcal{G}\ll T$, and let $t\in[0,T-\mathcal{G}]$. We define the two stopping times
@@
with the convention that $\inf\,\emptyset=\infty$. The stopping time $\tau^{+}_{t}$ is the first time after $t$ that the best bid price is above the best ask price at time $t$. If $\tau^{+}_{t}$$\infty$, a buy trade executed at $S_{t}^{a}$ becomes profitable for the client at time $\tau^{+}_{t}$ before the end of the trading day because the client can unwind her position and collect the profit
@@
Let $\mathcal{G}>0$ be a toxicity horizon. A client’s buy (resp. sell) filled by the broker at time $t$ is toxic for the broker if $\tau^{+}_{t}\leq t+\mathcal{G}$ (resp. if $\tau^{-}_{t}\leq t+\mathcal{G}$).
The above definition captures the broker’s exposure to adverse selection. A trade is labelled as toxic if over a given time window the client had the option to unwind the trade at a profit, in which case it would be a loss-leading trade for the broker. However, one cannot verify, ultimately, if the potentially toxic trade materialised as a loss to the broker or to another market participant. We do not have enough information to track each step, or potential step, in the life cycle of a trade to determine who made a loss or a gain — to make this assessment requires perfect knowledge of all trades by all market participants. [^4] There are alternative ways of defining toxic flow with our proprietary data set. For instance, the broker values a trade after unwinding it with the first trade in the opposite direction, i.e., a first-in-first-out inventory valuation.
@@
Figure 2: A client’s sell trade that becomes toxic for the broker after a few seconds of filling the trade. The $x$-axis is time and the $y$-axis is in units of USD.
In the figure, if a client buys from the broker at the best ask price at time t=t=10:00:00 am, then there is no opportunity for the trader to unwind the trade at a profit in the first ten seconds after the trade. However, had the trader sold to the broker at the best bid price at time t=t=10:00:00 am, then shortly after 10:00:01 the trade would be in-the-money for the client (i.e., toxic for the broker).
3.1 Toxicity profiles
@@
toxicity horizon $\mathcal{G}$ in seconds
@@
As expected, for short toxicity horizons only a small proportion of trades are toxic (e.g., 6.6% for a one second horizon), but as the toxicity horizon increases, the proportion of toxic trades grows considerably (e.g., it is roughly 70% after one minute). A simple mathematical argument can help us justify what we observe in the data. Consider a trader who sends a liquidity taking trade to the broker when the spread in the market is $\mathfrak{s}$ and suppose that the profitability of unwinding the trade, which is $\mathfrak{s}$ at time zero, diffuses according to a scaled Brownian motion $\sigma W_{t}$ with $\sigma>0$. From the reflection principle, the probability that such a trade becomes toxic at any point between zero and $\mathcal{G}$ seconds is given by
$\mathbb{P}\bigg(\sup_{t\in[0,\mathcal{G}]}\sigma\,W_{t}\geq\mathfrak{s}\bigg)=\mathbb{P}\bigg(\sup_{t\in[0,\mathcal{G}]}W_{t}\geq\frac{\mathfrak{s}}{\sigma}\bigg)=2\,\mathbb{P}\bigg(W_{\mathcal{G}}\geq\frac{\mathfrak{s}}{\sigma}\bigg)=2\,\bigg(1-\Phi\Big(\frac{\mathfrak{s}}{\sigma\sqrt{\mathcal{G}}}\Big)\bigg)\,$, (3.3)
where $\Phi$ is the standard normal cumulative distribution function. As the horizon $\mathcal{G}\to\infty$, the probability that the trade is toxic at some point converges to one. This does not mean that the broker loses money on each trade. That would happen only if the broker failed to hedge in the lit market and every liquidity unwound exited at the first profitable moment. In reality, the broker’s inventory shifts constantly due to both incoming trades and her own hedging in the lit market.
Arguably, a client can be labelled as toxic according to the percentage of their trades that are in-the-money after a set time frame, e.g., after $\mathcal{G}$ in Definition 1 above. Figure 3 shows the toxicity profiles of two clients trading EUR/USD on 8 July 2022 with LMAX Broker. The figures summarise all trades by clients A and B as follows. For each trade on 8 July 2022, the plot shows the profitability, from the client’s perspective, of unwinding each trade a given number of seconds after the trade. Here, the $x$-axis goes from zero to ten seconds and the $y$-axis is in dollars per million euros traded.
Figure 3: Profitability in dollars per million euros traded after a trade is executed. Panels correspond to two different clients. Blue line is the median trajectory and grey region is the 90% trajectory region. The $x$-axis is time and the $y$-axis is the profitability from the point of view of the client.
From Figure 3, and all else being equal, a broker would prefer to provide liquidity to client B instead of client A. More than 50% of the times that client A trades, the broker is exposed to making a loss on the trade in less than half a second. On the other hand, the median trajectory of profitability (blue line) for client B is below zero. [^5] Clearly, if the broker uses the flow from client A to inform other decisions and other investment strategies that are profitable for the broker, the net effect of trading at a loss with client A could be positive for the broker; see Cartea and Sánchez-Betancourt (2025).
@@ 3.2 Features to predict toxic trades
For each client, the broker uses features that reflect (i) the state of the LOB, (ii) recent activity in the LOB, and (iii) the cash and inventory of the client in the EUR/USD currency pair. [^6] We use the transactions with the broker to compute the cash and inventory of the clients. Clients can trade elsewhere but this is unknown to the broker. Our features build on Aït-Sahalia et al. (2022), who propose using three different clocks to aggregate LOB data. Here, we compute eight LOB statistics, each with three clocks (time, volume, and transaction), and seven backward-looking clock intervals of increasing length; thus, we have $8\times 3\times 7=168$ clock-based features. We also employ fifteen other features (e.g., cash and inventory), so there are 183 features per client.
@@
For each clock and a given interval in the past (as measured by the clock), we employ the following eight features: (a) volatility of the midprice, (b) number of trades executed by the client in the interval, (c) number of updates in the best quotes of LMAX Exchange in the interval, (d) return of the midprice over the interval, (e) average transformed volume in the best bid price, (f) average transformed volume in the best ask price, (g) average spread, and (h) average imbalance of the best available volumes, where “average” is simple arithmetic mean within an interval.
The three clocks provide alternative ways of deciding the datapoints that fall in a time interval. For example, with a time clock, the spread over the last two seconds is computed with all datapoints in the last 2 seconds. With a volume clock, the spread over the last $V$ units of volume traded is computing with all the data in the past (chronologically) until we gather $V$ units of volume traded. With a transaction clock, the spread over the last $k$ transactions is computed with the datapoints from the last $k$ transactions. See A for more details.